# Intuitionistic NF set theory

### v5

@@ 7. Cardinal exponentiation
$SSC(\{{\sf one},{\sf two}\})=\{\Lambda,\{{\sf one}\},\{{\sf two}\},\{{\sf one},{\sf two}\}\}\in 2^{\sf two}$
One can explicitly exhibit the ordered pairs of a similarity between the last-mentioned set and $\{{\sf one},{\sf two},{\sf three},{\sf four}\}\in{\sf four}$. We omit the details. Then by Lemma 4.8, $2^{\sf two}={\sf four}$. That completes the proof of the lemma.
@@
Proof. We have ${\sf two}={\sf one}^{+}$. If $a\neq b$ then by Lemma 6.5, $\{a\}\in{\sf one}$, and $\{a\}\cup\{b\}=\{a,b\}\in{\sf two}$. Conversely, If $u\in{\sf two}$ then $u=v\cup\{b\}$, where $v\in{\sf one}$ and $b\not\in v$. By Lemma 6.5, $v=\{a\}$ for some $a$, so $u=\{a,b\}$. That completes the proof of the lemma.
@@
Proof. We have ${\sf three}={\sf two}^{+}$. Assume $a,b,c$ are pairwise distinct. Then by Lemma 7.9, $\{a,b\}\in{\sf two}$. Since ${\sf three}=two^{+}$, $\{a,b\}\cup\{c\}=\{a,b,c\}\in{\sf three}$. Conversely, If $u\in{\sf three}$ then $u=v\cup\{c\}$, where $v\in{\sf two}$ and $c\not\in v$. By Lemma 7.9, $v=\{a,b\}$ for some $a,b$ with $a\neq b$. Since $c\not\in v$, $a\neq c$ and $b\neq c$. Therefore $u=\{a,b,c\}$ with $a,b,c$ pairwise distinct. That completes the proof of the lemma.
@@
We have ${\sf zero}<{\sf one}<{\sf two}<{\sf three}<four$.
Proof. Since each of these numbers is defined as the successor of the one listed just before it, the lemma is a consequence of Lemma 5.36.
@@
Proof. Let $m\in{\mathbb{F}}$ and $m<{\sf one}$. By Theorem 5.17, $m<{\sf zero}\ \lor\ ={\sf zero}\ \lor\ {\sf zero}<m$. By Lemma 5.28, $m<{\sf zero}$ is ruled out. It remains to rule out ${\sf zero}<m$. Assume ${\sf zero}<m$. Then
@@
Proof. Left to right. Assume $m\leq{\sf two}$. We have
@@
$a\in m\land a\in SSC(\{{\sf zero},{\sf one}\})$   for some $a$, by Lemma 5.9
@@
${\sf zero}\in a\lor{\sf zero}\not\in a$   since $a\in SSC(\{{\sf zero},{\sf one}\})$
${\sf one}\in a\lor{\sf one}\not\in a$   since $a\in SSC(\{{\sf zero},{\sf one}\})$
An argument by cases (about 170 steps, which we omit) shows that $a=\Lambda$, or $a=\{{\sf zero}\}$, or $a=\{{\sf one}\}$. Then $a={\sf zero}$ or ${\sf one}$, by Lemma 5.7. That completes the left to right direction.
Right to left: we have ${\sf zero}<{\sf two}$ and ${\sf one}<{\sf two}$ by Lemma 7.11. That completes the proof of the lemma.
@@
For all $a$, $a$ is a unit class if and only if $USC(a)$ is a unit class.
Proof. Left to right. Suppose $a=\{x\}$. Then the only unit subset of $a$ is $\{a\}$, so $USC(a)$ is a unit class.
Right to left. Suppose $USC(a)=\{u\}$. Then $u\in a$. Let $t\in a$. Then $\{t\}\in USC(a)$, so $\{t\}=u$. Hence every element of $a$ is equal to $u$. Hence $a=\{u\}$. That completes the proof of the lemma.
@@
Proof. Suppose $x\in{\mathbb{F}}$ and $x\leq{\sf zero}$. By the definition of $\leq$, there exists $a,b$ such that $a\in x$, $b\in{\sf zero}$, $a\subseteq b$, and $b=(a\cup b)-a$. Then
$b=\Lambda$   by definition of ${\sf zero}$
$a=\Lambda$   since $a\subseteq b$
$\Lambda\in x\cap{\sf zero}$   by definition of $\cap$
@@
That completes the proof of the lemma.
Lemma 7.16 (Specker 4.6*).
@@
Proof. Since ${\mathbb{F}}$ has decidable equality, by Corollary 5.18
@@
Case 1, $m={\sf zero}$. Then by Lemma 7.6, $2^{m}={\sf one}$, and we have to show ${\sf zero}<{\sf one}$, which follows from the definition of $<$ by exhibiting the separable subset $\Lambda$ of the set $\{\Lambda\}$, and noting that $\Lambda\in{\sf zero}$ while $\{\Lambda\}\in{\sf one}$.
Case 2, $m={\sf one}$. Then by Lemma 7.7, $2^{m}={\sf two}$, and we have to show ${\sf one}<{\sf two}$, which follows from ${\sf zero}<{\sf one}$ by Corollary 4.7 and Lemma 5.11, or more directly, from the definition of $<$ by exhibiting the separable subset $\{\Lambda\}$ of $\{\{\Lambda\},\Lambda\}$, the former of which belongs to ${\sf one}$ while the latter belongs to ${\sf two}$.
Case 3, $m\neq{\sf zero}$ and $m\neq{\sf one}$. By hypothesis, $2^{m}$ is inhabited. Then there exists $a$ such that $USC(a)\in m$. Since $m\in{\mathbb{F}}$, we have
$USC(a)\in{\sf FINITE}$   by Lemma 4.4
@@
Then by Lemma 3.18, $USC(a)$ is a separable subset of $SSC(a)$. Now with $u=USC(a)$ and $v=SSC(a)$ we have proved that $u$ is a separable subset of $v$ and $u\in m$ and $v\in 2^{m}$. Then by Definition 5.2 we have $m\leq 2^{m}$.
By definition $m<2^{m}$ means $m\leq 2^{m}$ and $m\neq 2^{m}$. It remains to prove that $m\neq 2^{m}$. Suppose $m=2^{m}$. As just proved, we have $USC(a)\subseteq SSC(a)$. I say that it is a proper subset, $USC(a)\subset SSC(a)$. It suffices to prove $USC(a)\neq SSC(a)$. We have to produce an element of $SSC(a)$ that does not belong to $USC(a)$. We propose $a$ as this element. We have $a\in SSC(a)$ since $a$ is a separable subset of itself. It remains to show that $a\not\in USC(a)$. Assume $a\in USC(a)$. Then $a$ is a unit class. By Lemma 7.14, $USC(a)$ is also a unit class. Any two unit classes are similar, so $USC(a)\sim{\sf zero}$. Since ${\sf zero}\in{\sf one}$, $USC(a)\in{\sf one}$, by Lemma 4.8. Then $m\cap{\sf one}$ is inhabited, since it contains $USC(a)$. Then by Lemma 5.7, $m={\sf one}$, contradiction. That completes the proof that $USC(a)$ is a proper subset of $SSC(a)$.
@@
$USC(a)\subset SSC(a)$   as proved above
$USC(a)\sim SSC(a)$   by Lemma 4.9, since $USC(a)\in m$ and $SSC(a)\in 2^{m}$
$SSC(a)\mbox{\ is infinite}$   since $SSC(a)\sim USC(a)\subset SSC(a)$
$\neg\,(SSC(a)\in{\sf FINITE})$   by Theorem 3.24
$SSC(a)\in{\sf FINITE}$   by Lemma 3.17, since $a\in{\sf FINITE}$
That is a contradiction. That completes the proof of the lemma.
@@
Proof. Suppose $m\in{\mathbb{F}}$ and $\exists u\,(u\in 2^{m})$. Then
@@
Proof. Suppose $m\leq n$ and $2^{n}$ is inhabited. Then
$\exists u\,(u\in n)$   by Corollary  4.7
$\exists b\,(USC(b)\in n)$   by Lemma 7.4
$\exists b\,(USC(b)\in m)$   by Lemma 7.4
Since $m\leq n$, by Lemma 5.9 there is a separable subset $x$ of $USC(b)$ such that $x\in m$. Let $a=\bigcup x$. Then using the definitions of $\bigcup$ and $USC$, we have $x=USC(a)$. Therefore $2^{m}$ is inhabited. Now $2^{m}=NC(SSC(a))$ and $2^{n}=NC(SSC(b))$.
@@
$USC(b)\in n$
$USC(b)\in{\sf FINITE}$   by Lemma 4.4
@@
Every member of $x$ is a unit class, since $x=USC(a)$). Every unit class is finite. Therefore every member of $x$ is finite. Moreover, since the members of $x$ are unit classes, distinct members of $x$ are disjoint. Since $x$ is also finite, $a=\bigcup x$ is a finite union of disjoint finite sets. Hence $a$ is finite, by Lemma 3.25.
Since $x=USC(a)$ is a separable subset of $USC(b)$, we have
$USC(a)\in SSC(USC(b))$
$a\in SSC(b)$   by Lemma 6.11 (right to left)
$SSC(a)\in SSC(SSC(b))$   by Lemma 6.14, since $a$ and $b$ are finite
Then $SSC(a)$ belongs to $2^{m}$, and is a separable subset of $SSC(b)$, which belongs to $2^{n}$. Therefore, by Definition 5.2, $2^{m}\leq 2^{n}$. That completes the proof of the lemma.
8. Addition
@@
$x+y:=\{z:\exists u,v\,(u\in x\ \land\ v\in y\ \land\ u\cap v=\Lambda\ \land\ z=u\cup v)\}$
The formula in the definition is stratified, giving $u$, $v$, and $z$ index 1 and $x$ and $y$ index 2. Then $x,y$, and $z$ all get the same index, so addition is definable as a function in INF. (See Definition 2.2 for ordered triples.)
@@
Proof. Ad $x+{\sf zero}=x$. By extensionality, it suffices to show $z\in x+{\sf zero}{\ \leftrightarrow\ }z\in x$. Left to right: suppose $z\in x+{\sf zero}$. Then $z=u\cup v$, where $u$ and $v$ are disjoint and $u\in x$ and $v\in{\sf zero}$. Since ${\sf zero}=\{\Lambda\}$, we have $v=\lambda$, so $z=u\cup\Lambda=u\in x$. That completes the left-to-right implication.
Right-to-left: Let $z\in x$. Then $z\cup\Lambda\in x+{\sf zero}$, by the definition of addition. Since $z\cup\Lambda=z$, we have $z\in x+{\sf zero}$ as desired. That completes the proof of $x+{\sf zero}=x$.
@@
$u\cap(v\cup\{c\})=\Lambda$   since $u\cap v=\Lambda$ and $c\not\in u$
@@
$(u\cup\{c\})\cap v=\Lambda$   by the associativity and commutativity of union
@@
That completes the right to left direction. That completes the proof of the lemma.
@@
Remark. We call attention to the fact that, even when $x,y,z$ are assumed to be in ${\mathbb{F}}$, the expressions in the equations might “overflow”, and the equations contain implicitly the assertion that the overflows “match”, i.e., one side overflows if and only if the other does. Here “overflow” means to have the value $\Lambda$.
Proof. These laws are immediate consequences of the definition of addition, via the associative and commutative laws of set union. We omit the proofs.
@@
Proof. By induction on $\mu$, which is legal since the formula is stratified.
@@
Induction step: Suppose $\kappa+\mu^{+}$ is inhabited and $\mu^{+}$ is inhabited. Then $\kappa+\mu^{+}=(\kappa+\mu)^{+}$ is inhabited. By the induction hypothesis, $\kappa+\mu\in{\mathbb{F}}$. Then by Lemma 4.19, $(\kappa+\mu)^{+}\in{\mathbb{F}}$. Since $(\kappa+\mu)^{+}=\kappa+\mu^{+}$, we have $\kappa+\mu^{+}\in{\mathbb{F}}$. That completes the induction step. That completes the proof of the lemma.
@@
Proof. By Corollary 4.7, $p+q+r$ is inhabited. Let $u\in p+q+r$. Then by the definition of addition, $u=a\cup b\cup c$ with $a\in p$, $b\in q$, $c\in r$, and $a,b,c$ pairwise disjoint. Then $a\cup b\in p+q$ and $b\cup c\in q+r$. Then by Lemma 8.6, $p+q\in{\mathbb{F}}$ and $q+r\in{\mathbb{F}}$. That completes the proof of the three summand case. The case of four summands is treated similarly. We omit the details.
@@
Proof. Suppose $p\in{\mathbb{F}}$ and $p+q^{+}\in{\mathbb{F}}$. By Corollary 4.7, there exists $u\in p+q^{+}$. Then by Definition 8.1, there exist $a$ and $b$ with $a\in p$ and $b\in q^{+}$ and $a\cap b=\Lambda$. By definition of successor, $b=x\cup\{c\}$ for some $x$ and $c$, so $c\in b$. Since $a\cap b=\Lambda$, we have $c\not\in a$. Then $a\cup\{c\}\in p^{+}$. Then $p^{+}\in{\mathbb{F}}$. That completes the proof of the lemma.
@@
Proof. Suppose $a,b,p,q\in{\mathbb{F}}$ and $b+q\in{\mathbb{F}}$. Suppose also $a\leq b$, $p\leq q$. Then
@@
By the definition of addition, there exist $u,v$ with $w=u\cup v$, $u\in b$, $v\in q$, and $u\cap v=\Lambda$. By Lemma 5.9, since $a\leq b$ there exists $r\in a$ with $r\in SSC(u)$. By Lemma 5.9, since $p\leq q$, there exists $s\in p$ with $s\in SSC(v)$. Then one can verify that $r\cup s\in SSC(u\cup v)$. (We omit the details of that verification.) Since $u\cup v=w$ we have $r\cup s\in SSC(w)$. We have $r\cap s=\Lambda$, since $r\subseteq u$, $s\subseteq v$, and $u\cap v=\Lambda$. Then $r\cup s\in a+p$, by the definition of addition. Then $a+p\leq b+q$, as witnessed by $r\cup s\in a+p$, $r\cup s\in SSC(w)$, and $w\in b+q$. That completes the proof of the first assertion of the lemma.
@@
Proof. Suppose $a<b$ and $p\leq q$. Then
@@
$USC(x)\in m\ {\rightarrow}\ SSC(x)\in 2^{m}.$
Proof. Suppose $USC(x)\in m$. By Definition 7.1, $2^{m}$ contains all sets similar to $SSC(x)$. By Lemma 2.11, $SSC(x)$ is one of those sets, so $SSC(x)\in 2^{m}$. That completes the proof of the lemma.
@@
Proof. Suppose $2^{z}={\sf zero}$. Then
$\Lambda\in{\sf zero}$   by Definition 5.15
$\Lambda\in 2^{z}$   since $2^{z}={\sf zero}$
$\Lambda~{}\sim SSC(a)\ \land\ USC(a)\in x$   by Definition 7.1
$SSC(a)=\Lambda$   since only $\Lambda$ is similar to $\Lambda$
But $a\in SSC(a)$, contradiction. That completes the proof of the lemma.
@@
Proof. Extend a similarity $f:x\to y$ by defining $f(a)=b$. We omit the details.
@@
Let $p$ and $q$ be disjoint finite sets. Then $Nc(p\cup q)=Nc(p)+Nc(q)$.
Proof. We have
@@
$Nc(p\cup q)\in{\mathbb{F}}$   by Lemma 4.21
$Nc(p)\in{\mathbb{F}}$   by Lemma 4.21
$Nc(q)\in{\mathbb{F}}$   by Lemma 4.21
$p\cup q\in Nc(p\cup q)$   by Lemma 4.11
$p\in Nc(p)$   by Lemma 4.11
$q\in Nc(q)$   by Lemma 4.11
$p\cup q\in Nc(p)+Nc(q)$   by the definition of addition
$Nc(p\cup q)\cap Nc(p)+Nc(q)\neq\Lambda$   since both contain $p\cup q$
$Nc(p)+Nc(q)\in{\mathbb{F}}$   by Lemma 8.6
$Nc(p\cup q)=Nc(p)+Nc(q)$   by Lemma 5.7
That completes the proof of the lemma.
@@
Proof. The two formulas are equivalent, by Lemma 8.3. We prove the first one by induction on $p$, which is legal since the formula is stratified. More precisely we prove by induction on $p$ that
@@
That completes the induction step. That completes the proof of the lemma.

### v6

@@ 7 Cardinal exponentiation
$\mathcal{P}_{s}(\{{\sf one},{\sf two}\})=\{\emptyset,\{{\sf one}\},\{{\sf two}\},\{{\sf one},{\sf two}\}\}\in 2^{\sf two}$
One can explicitly exhibit the ordered pairs of a similarity between the last-mentioned set and the element $\{{\sf one},{\sf two},{\sf three},{\sf four}\}$ of $\sf four$}. We omit the details. Then by Lemma 4.8, $2^{\sf two}={\sf four}$. ∎
@@
Proof.
We have ${\sf two}={\sf one}^{+}$. If $a\neq b$ then by Lemma 6.5, $\{a\}\in{\sf one}$, and $\{a\}\cup\{b\}=\{a,b\}\in{\sf two}$. Conversely, If $u\in{\sf two}$ then $u=v\cup\{b\}$, where $v\in{\sf one}$ and $b\not\in v$. By Lemma 6.5, $v=\{a\}$ for some $a$, so $u=\{a,b\}$. ∎
@@
Proof.
We have ${\sf three}={\sf two}^{+}$. Assume $a,b,c$ are pairwise distinct. Then by Lemma 7.9, $\{a,b\}\in{\sf two}$. Since ${\sf three}=two^{+}$, $\{a,b\}\cup\{c\}=\{a,b,c\}\in{\sf three}$. Conversely, If $u\in{\sf three}$ then $u=v\cup\{c\}$, where $v\in{\sf two}$ and $c\not\in v$. By Lemma 7.9, $v=\{a,b\}$ for some $a,b$ with $a\neq b$. Since $c\not\in v$, $a\neq c$ and $b\neq c$. Therefore $u=\{a,b,c\}$ with $a,b,c$ pairwise distinct. ∎
@@
We have ${\sf zero}<{\sf one}<{\sf two}<{\sf three}<{\sf four}$.
Proof.
Since each of these numbers is defined as the successor of the one listed just before it, the lemma is a consequence of Lemma 5.37. ∎
@@
Proof.
Let $m\in{\mathbb{F}}$ and $m<{\sf one}$. By Theorem 5.16, $m<{\sf zero}\ \lor\ ={\sf zero}\ \lor\ {\sf zero}<m$. By Lemma 5.29, $m<{\sf zero}$ is ruled out. It remains to rule out ${\sf zero}<m$. Assume ${\sf zero}<m$. Then
@@
Proof.
Left to right. Assume $m\leq{\sf two}$. We have
@@
$a\in m\land a\in\mathcal{P}_{s}(\{{\sf zero},{\sf one}\})$   for some $a$, by Lemma 5.8
@@
${\sf zero}\in a\lor{\sf zero}\not\in a$   since $a\in\mathcal{P}_{s}(\{{\sf zero},{\sf one}\})$
${\sf one}\in a\lor{\sf one}\not\in a$   since $a\in\mathcal{P}_{s}(\{{\sf zero},{\sf one}\})$
An argument by cases (about 170 steps, which we omit) shows that $a=\emptyset$, or $a=\{{\sf zero}\}$, or $a=\{{\sf one}\}$. Then $a={\sf zero}$ or $\sf one$}, by Lemma 4.24. That completes the left to right direction.
Right to left: we have ${\sf zero}<{\sf two}$ and ${\sf one}<{\sf two}$ by Lemma 7.11. ∎
@@
For all $a$, $a$ is a unit class if and only if $\mathcal{P}_{1}(a)$ is a unit class.
Proof.
Left to right. Suppose $a=\{x\}$. Then the only unit subset of $a$ is $\{a\}$, so $\mathcal{P}_{1}(a)$ is a unit class.
Right to left. Suppose $\mathcal{P}_{1}(a)=\{u\}$. Then $u\in a$. Let $t\in a$. Then $\{t\}\in\mathcal{P}_{1}(a)$, so $\{t\}=u$. Hence every element of $a$ is equal to $u$. Hence $a=\{u\}$. ∎
@@
Proof.
Suppose $x\in{\mathbb{F}}$ and $x\leq{\sf zero}$. By the definition of $\leq$, there exists $a,b$ such that $a\in x$, $b\in{\sf zero}$, $a\subseteq b$, and $b=(a\cup b)-a$. Then
$b=\emptyset$   by definition of $\sf zero$}
$a=\emptyset$   since $a\subseteq b$
$\emptyset\in x\cap{\sf zero}$   by definition of $\cap$
@@
Lemma 7.16 (Specker 4.6).
@@
Remark. This version of Specker 4.6 phrases the matter positively, so it is constructively stronger.
Proof.
Since $\mathbb{F}$} has decidable equality, by Corollary 5.17
@@
Case 1, $m={\sf zero}$. Then by Lemma 7.6, $2^{m}={\sf one}$, and we have to show ${\sf zero}<{\sf one}$, which follows from the definition of << by exhibiting the separable subset $\emptyset$ of the set $\{\emptyset\}$, and noting that $\emptyset\in{\sf zero}$ while $\{\emptyset\}\in{\sf one}$.
Case 2, $m={\sf one}$. Then by Lemma 7.7, $2^{m}={\sf two}$, and we have to show ${\sf one}<{\sf two}$, which follows from ${\sf zero}<{\sf one}$ by Corollary 4.7 and Lemma 5.10, or more directly, from the definition of << by exhibiting the separable subset $\{\emptyset\}$ of $\{\{\emptyset\},\emptyset\}$, the former of which belongs to $\sf one$} while the latter belongs to $\sf two$}.
Case 3, $m\neq{\sf zero}$ and $m\neq{\sf one}$. By hypothesis, $2^{m}$ is inhabited. Then there exists $a$ such that $\mathcal{P}_{1}(a)\in m$. Since $m\in{\mathbb{F}}$, we have
$\mathcal{P}_{1}(a)\in{\sf FINITE}$   by Lemma 4.4
@@
Then by Lemma 3.18, $\mathcal{P}_{1}(a)$ is a separable subset of $\mathcal{P}_{s}(a)$. Now with $u=\mathcal{P}_{1}(a)$ and $v=\mathcal{P}_{s}(a)$ we have proved that $u$ is a separable subset of $v$ and $u\in m$ and $v\in 2^{m}$. Then by Definition 5.2 we have $m\leq 2^{m}$.
By definition $m<2^{m}$ means $m\leq 2^{m}$ and $m\neq 2^{m}$. It remains to prove that $m\neq 2^{m}$. Suppose $m=2^{m}$. As just proved, we have $\mathcal{P}_{1}(a)\subseteq\mathcal{P}_{s}(a)$. I say that it is a proper subset, $\mathcal{P}_{1}(a)\subset\mathcal{P}_{s}(a)$. It suffices to prove $\mathcal{P}_{1}(a)\neq\mathcal{P}_{s}(a)$. We have to produce an element of $\mathcal{P}_{s}(a)$ that does not belong to $\mathcal{P}_{1}(a)$. We propose $a$ as this element. We have $a\in\mathcal{P}_{s}(a)$ since $a$ is a separable subset of itself. It remains to show that $a\not\in\mathcal{P}_{1}(a)$. Assume $a\in\mathcal{P}_{1}(a)$. Then $a$ is a unit class. By Lemma 7.14, $\mathcal{P}_{1}(a)$ is also a unit class. Any two unit classes are similar, so $\mathcal{P}_{1}(a)\sim{\sf zero}$. Since ${\sf zero}\in{\sf one}$, $\mathcal{P}_{1}(a)\in{\sf one}$, by Lemma 4.8. Then $m\cap{\sf one}$ is inhabited, since it contains $\mathcal{P}_{1}(a)$. Then by Lemma 4.24, $m={\sf one}$, contradiction. That completes the proof that $\mathcal{P}_{1}(a)$ is a proper subset of $\mathcal{P}_{s}(a)$.
@@
$\mathcal{P}_{1}(a)\subset\mathcal{P}_{s}(a)$   as proved above
$\mathcal{P}_{1}(a)\sim\mathcal{P}_{s}(a)$   by Lemma 4.9, since $\mathcal{P}_{1}(a)\in m$ and $\mathcal{P}_{s}(a)\in 2^{m}$
$\mathcal{P}_{s}(a)\mbox{\ is infinite}$   since $\mathcal{P}_{s}(a)\sim\mathcal{P}_{1}(a)\subset\mathcal{P}_{s}(a)$
$\neg\,(\mathcal{P}_{s}(a)\in{\sf FINITE})$   by Theorem 3.24
$\mathcal{P}_{s}(a)\in{\sf FINITE}$   by Lemma 3.17, since $a\in{\sf FINITE}$
That is a contradiction. ∎
@@
Proof.
Suppose $m\in{\mathbb{F}}$ and $\exists u\,(u\in 2^{m})$. Then
@@
Proof.
Suppose $m\leq n$ and $2^{n}$ is inhabited. Then
$\exists u\,(u\in n)$   by Corollary 4.7
$\exists b\,(\mathcal{P}_{1}(b)\in n)$   by Lemma 7.4
$\exists b\,(\mathcal{P}_{1}(b)\in m)$   by Lemma 7.4
Since $m\leq n$, by Lemma 5.8 there is a separable subset $x$ of $\mathcal{P}_{1}(b)$ such that $x\in m$. Let $a=\bigcup x$. Then using the definitions of $\bigcup$ and $\mathcal{P}_{1}$, we have $x=\mathcal{P}_{1}(a)$. Therefore $2^{m}$ is inhabited. Now $2^{m}=NC(\mathcal{P}_{s}(a))$ and $2^{n}=NC(\mathcal{P}_{s}(b))$.
@@
$\mathcal{P}_{1}(b)\in n$
$\mathcal{P}_{1}(b)\in{\sf FINITE}$   by Lemma 4.4
@@
Every member of $x$ is a unit class, since $x=\mathcal{P}_{1}(a)$). Every unit class is finite. Therefore every member of $x$ is finite. Moreover, since the members of $x$ are unit classes, distinct members of $x$ are disjoint. Since $x$ is also finite, $a=\bigcup x$ is a finite union of disjoint finite sets. Hence $a$ is finite, by Lemma 3.25.
Since $x=\mathcal{P}_{1}(a)$ is a separable subset of $\mathcal{P}_{1}(b)$, we have
$\mathcal{P}_{1}(a)\in\mathcal{P}_{s}(\mathcal{P}_{1}(b))$
$a\in\mathcal{P}_{s}(b)$   by Lemma 6.11 (right to left)
$\mathcal{P}_{s}(a)\in\mathcal{P}_{s}(\mathcal{P}_{s}(b))$   by Lemma 6.14, since $a$ and $b$ are finite
Then $\mathcal{P}_{s}(a)$ belongs to $2^{m}$, and is a separable subset of $\mathcal{P}_{s}(b)$, which belongs to $2^{n}$. Therefore, by Definition 5.2, $2^{m}\leq 2^{n}$. ∎
8 Addition
@@
$x+y:=\{z:\exists u,v\,(u\in x\ \land\ v\in y\ \land\ u\cap v=\emptyset\ \land\ z=u\cup v)\}$
The formula in the definition is stratified, giving $u$, $v$, and $z$ index 1 and $x$ and $y$ index 2. Then $x,y$, and $z$ all get the same index, so addition is definable as a function in $\mathit{i}\mkern 0.4mu\mathrm{NF}$. (See Definition 2.2 for ordered triples.)
@@
Proof.
Ad $x+{\sf zero}=x$. By extensionality, it suffices to show $z\in x+{\sf zero}{\ \leftrightarrow\ }z\in x$. Left to right: suppose $z\in x+{\sf zero}$. Then $z=u\cup v$, where $u$ and $v$ are disjoint and $u\in x$ and $v\in{\sf zero}$. Since $\sf zero$$\{\emptyset\}$, we have $v=\lambda$, so $z=u\cup\emptyset=u\in x$. That completes the left-to-right implication.
Right-to-left: Let $z\in x$. Then $z\cup\emptyset\in x+{\sf zero}$, by the definition of addition. Since $z\cup\emptyset=z$, we have $z\in x+{\sf zero}$ as desired. That completes the proof of $x+{\sf zero}=x$.
@@
$u\cap(v\cup\{c\})=\emptyset$   since $u\cap v=\emptyset$ and $c\not\in u$
@@
$(u\cup\{c\})\cap v=\emptyset$   by the associativity and commutativity of union
@@
That completes the right to left direction. ∎
@@
Remark. We call attention to the fact that, even when $x,y,z$ are assumed to be in $\mathbb{F}$}, the expressions in the equations might “overflow”, and the equations contain implicitly the assertion that the overflows “match”, i.e., one side overflows if and only if the other does. Here “overflow” means to have the value $\emptyset$.
Proof.
These laws are immediate consequences of the definition of addition, via the associative and commutative laws of set union. We omit the proofs. ∎
@@
Proof.
By induction on $\mu$, which is legal since the formula is stratified.
@@
Induction step: Suppose $\kappa+\mu^{+}$ is inhabited and $\mu^{+}$ is inhabited. Then $\kappa+\mu^{+}=(\kappa+\mu)^{+}$ is inhabited. By the induction hypothesis, $\kappa+\mu\in{\mathbb{F}}$. Then by Lemma 4.19, $(\kappa+\mu)^{+}\in{\mathbb{F}}$. Since $(\kappa+\mu)^{+}=\kappa+\mu^{+}$, we have $\kappa+\mu^{+}\in{\mathbb{F}}$. That completes the induction step. ∎
@@
Proof.
By Corollary 4.7, $p+q+r$ is inhabited. Let $u\in p+q+r$. Then by the definition of addition, $u=a\cup b\cup c$ with $a\in p$, $b\in q$, $c\in r$, and $a,b,c$ pairwise disjoint. Then $a\cup b\in p+q$ and $b\cup c\in q+r$. Then by Lemma 8.6, $p+q\in{\mathbb{F}}$ and $q+r\in{\mathbb{F}}$. That completes the proof of the three summand case. The case of four summands is treated similarly. We omit the details. ∎
@@
Proof.
Suppose $p\in{\mathbb{F}}$ and $p+q^{+}\in{\mathbb{F}}$. By Corollary 4.7, there exists $u\in p+q^{+}$. Then by Definition 8.1, there exist $a$ and $b$ with $a\in p$ and $b\in q^{+}$ and $a\cap b=\emptyset$. By definition of successor, $b=x\cup\{c\}$ for some $x$ and $c$, so $c\in b$. Since $a\cap b=\emptyset$, we have $c\not\in a$. Then $a\cup\{c\}\in p^{+}$. Then $p^{+}\in{\mathbb{F}}$. ∎
@@
Proof.
Suppose $a,b,p,q\in{\mathbb{F}}$ and $b+q\in{\mathbb{F}}$. Suppose also $a\leq b$, $p\leq q$. Then
@@
By the definition of addition, there exist $u,v$ with $w=u\cup v$, $u\in b$, $v\in q$, and $u\cap v=\emptyset$. By Lemma 5.8, since $a\leq b$ there exists $r\in a$ with $r\in\mathcal{P}_{s}(u)$. By Lemma 5.8, since $p\leq q$, there exists $s\in p$ with $s\in\mathcal{P}_{s}(v)$. Then one can verify that $r\cup s\in\mathcal{P}_{s}(u\cup v)$. (We omit the details of that verification.) Since $u\cup v=w$ we have $r\cup s\in\mathcal{P}_{s}(w)$. We have $r\cap s=\emptyset$, since $r\subseteq u$, $s\subseteq v$, and $u\cap v=\emptyset$. Then $r\cup s\in a+p$, by the definition of addition. Then $a+p\leq b+q$, as witnessed by $r\cup s\in a+p$, $r\cup s\in\mathcal{P}_{s}(w)$, and $w\in b+q$. ∎
@@
Proof.
Suppose $a<b$ and $p\leq q$. Then
@@
$\mathcal{P}_{1}(x)\in m\ {\rightarrow}\ \mathcal{P}_{s}(x)\in 2^{m}.$
Proof.
Suppose $\mathcal{P}_{1}(x)\in m$. By Definition 7.1, $2^{m}$ contains all sets similar to $\mathcal{P}_{s}(x)$. By Lemma 2.11, $\mathcal{P}_{s}(x)$ is one of those sets, so $\mathcal{P}_{s}(x)\in 2^{m}$. ∎
@@
Proof.
Suppose $2^{z}={\sf zero}$. Then
$\emptyset\in{\sf zero}$   by Definition 5.14
$\emptyset\in 2^{z}$   since $2^{z}={\sf zero}$
$\emptyset~\sim\mathcal{P}_{s}(a)\ \land\ \mathcal{P}_{1}(a)\in x$   by Definition 7.1
$\mathcal{P}_{s}(a)=\emptyset$   since only $\emptyset$ is similar to $\emptyset$
But $a\in\mathcal{P}_{s}(a)$, contradiction. ∎
@@
Proof.
Extend a similarity $f:x\to y$ by defining $f(a)=b$. We omit the details. ∎
@@
Let $p$ and $q$ be disjoint finite sets. Then $|p\cup q|=|p|+|q|$.
Proof.
We have
@@
$|p\cup q|\in{\mathbb{F}}$   by Lemma 4.21
$|p|\in{\mathbb{F}}$   by Lemma 4.21
$|q|\in{\mathbb{F}}$   by Lemma 4.21
$p\cup q\in|p\cup q|$   by Lemma 4.11
$p\in|p|$   by Lemma 4.11
$q\in|q|$   by Lemma 4.11
$p\cup q\in|p|+|q|$   by the definition of addition
$|p\cup q|\cap|p|+|q|\neq\emptyset$   since both contain $p\cup q$
$|p|+|q|\in{\mathbb{F}}$   by Lemma 8.6
$|p\cup q|=|p|+|q|$   by Lemma 4.24
@@
Proof.
The two formulas are equivalent, by Lemma 8.3. We prove the first one by induction on $p$, which is legal since the formula is stratified. More precisely we prove by induction on $p$ that
@@
That completes the induction step. ∎